# A generalization of the relaxation-based matrix splitting iterative method for solving the system of generalized absolute value equations

### v1

@@
Abstract: By incorporating a new matrix splitting and the momentum acceleration into the relaxed-based matrix splitting (RMS) method [26], a generalization of the RMS (GRMS) iterative method for solving the generalized absolute value equations (GAVEs) is proposed. Unlike some existing methods, by using the Cauchy’s convergence principle, we give some sufficient conditions for the existence and uniqueness of the solution to the GAVEs and prove that our method can converge to the unique solution of the GAVEs. Moreover, we can obtain a few new and weaker convergence conditions for some existing methods. Preliminary numerical experiments show that the proposed method is efficient.
Keyword: Generalized absolute value equations; Matrix splitting; Relaxation technique; Momentum acceleration; Convergence.
@@ 1 Introduction
In this paper, we focus on seeking the solution of the following system of generalized absolute value equations (GAVEs)
$Ax-B|x|-b=0,$ (1.1)
where $A,B\in\mathbb{R}^{n\times n}$ and $b\in\mathbb{R}^{n}$ are given, and $x\in\mathbb{R}^{n}$ is unknown. Here, $|x|\in\mathbb{R}^{n}$ denotes the vector whose $i$th component is $|x_{i}|$, i.e., the absolute value of $x_{i}$. When $B=I$, the GAVEs (1.1) reduces to the system of absolute value equations (AVEs)
$Ax-|x|-b=0.$ (1.2)
To our knowledge, GAVEs (1.1) is formally introduced by Rohn in [24]. Over the past two decades, GAVEs (1.1) and AVEs (1.2) have attracted more and more attention in the optimization community and a great deal of research has been done theoretically and numerically; see, e.g., [16, 6, 8, 12, 17, 18, 19, 3, 11, 1, 20, 4]. Particularly, solving the GAVEs (1.1) is NP-hard [17].
Within the numerical algorithms for solving the GAVEs (1.1), some are characterized by using the matrix splitting technique or the relaxation technique. Recently, by using the matrix splitting technique and the relaxation technique, Song and Song [26] present the relaxation-based matrix splitting (RMS) iterative method (see Algorithm 1.1 below) for solving the GAVEs (1.1).
Algorithm 1.1.
Assume that $A=M_{\rm RMS}-N_{\rm RMS}$ with $M_{\rm RMS}$ being nonsingular. Given the initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$, for $k=0,1,2,\cdots$ until the iterative sequence $\{(x^{(k)},y^{(k)})\}^{\infty}_{k=0}$ is convergent, compute
$\begin{cases}x^{(k+1)}=M_{\rm RMS}^{-1}(N_{\rm RMS}x^{(k)}+By^{(k)}+b),\\ y^{(k+1)}=(1-\tau_{\rm RMS})y^{(k)}+\tau_{\rm RMS}|x^{(k+1)}|,\end{cases}$ (1.3)
@@
The RMS iterative scheme (1.3) has a general form and it includes many existing iterative schemes as its special cases [26]. For instance,
@@
    Let $A=M_{\rm NMS}-N_{\rm NMS}$ and $\Omega_{\rm NMS}$ be a given matrix such that $M_{\rm NMS}+\Omega_{\rm NMS}$ is invertible. If $\tau_{\rm RMS}=1$ and $M_{\rm RMS}=M_{\rm NMS}+\Omega_{\rm NMS}$, the RMS iteration (1.3) with $y^{(0)}=|x^{(0)}|$ turns into the Newton-based matrix splitting (NMS) iteration [33]
    $x^{(k+1)}=(M_{\rm NMS}+\Omega_{\rm NMS})^{-1}[(N_{\rm NMS}+\Omega_{\rm NMS})x^{(k)}+B|x^{(k)}|+b].$ (1.4)
    The NMS iteration (1.4) reduces to the modified Newton-type (MN) iteration [27] when $M_{\rm NMS}=A$ and $\Omega_{\rm NMS}$ is a positive semi-definite matrix. Moreover, if $M_{\rm NMS}=A$ and $\Omega_{\rm NMS}=0$, the NMS iteration (1.4) reduces to the Picard iteration [25]
    $x^{(k+1)}=A^{-1}(B|x^{(k)}|+b).$ (1.5)
    If $M_{\rm NMS}=\dfrac{1}{2}(A+\Omega_{\rm SSMN})$ and $\Omega_{\rm NMS}=0$, then the NMS iteration (1.4) turns into the following shift splitting modified Newton-type (SSMN) iteration [14]
    $x^{(k+1)}=(A+\Omega_{\rm SSMN})^{-1}\left[(\Omega_{\rm SSMN}-A)x^{(k)}+2B|x^{(k)}|+2b\right],$ (1.6)
    where $\Omega_{\rm SSMN}\in\mathbb{R}^{n\times n}$ is a positive semi-definite matrix such that $A+\Omega_{\rm SSMN}$ is invertible.
@@
    If $B=I,M_{\rm RMS}=\dfrac{1}{\omega_{\rm SOR}}A$ and $\tau_{\rm RMS}=\omega_{\rm SOR}>0$, the RMS iteration (1.3) becomes the SOR-like iteration [9]
    $\begin{cases}x^{(k+1)}=(1-\omega_{\rm SOR})x^{(k)}+\omega_{\rm SOR}A^{-1}\left(y^{(k)}+b\right),\\ y^{(k+1)}=(1-\omega_{\rm SOR})y^{(k)}+\omega_{\rm SOR}|x^{(k+1)}|.\end{cases}$ (1.7)
@@
    Throughout this paper, let $A=D-L-U$ with $D$, $-L$ and $-U$ being the diagonal part, the strictly lower-triangular part and the strictly upper-triangular part of $A$, respectively. If $B=I$, $M_{\rm RMS}=\dfrac{1}{\omega}D-L$ and $\tau_{\rm RMS}=\omega>0$, then the RMS iteration (1.3) turns into the modified SOR-like iteration [13, Equation (11)].
@@
    If $B=I,M_{\rm RMS}=A,\tau_{\rm RMS}=\tau_{\rm FPI}>0$, then the RMS iteration (1.3) is reduced to the fixed point iteration (FPI) [10]
    $\begin{cases}x^{(k+1)}=A^{-1}(y^{(k)}+b),\\ y^{(k+1)}=\left(1-\tau_{\rm FPI}\right)y^{(k)}+\tau_{\rm FPI}|x^{(k+1)}|,\end{cases}$ (1.8)
@@
    $\begin{cases}x^{(k+1)}=x^{(k)}+\omega_{\rm NSOR}A^{-1}(b-Ax^{(k)}+y^{(k)}),\\ y^{(k+1)}=y^{(k)}+\dfrac{\omega_{\rm NSOR}}{\sigma_{\rm NSOR}}(b+|x^{(k+1)}|-Ax^{(k+1)}),\end{cases}$ (1.9)
[…]
Before going ahead, we should point out that the FPI (1.8) is also a special case of the modified FPI (MFPI) [29]
$\begin{cases}x^{(k+1)}=A^{-1}(Q_{\rm MFPI}y^{(k)}+b),\\ y^{(k+1)}=(1-\tau_{\rm MFPI})y^{(k)}+\tau_{\rm MFPI}Q^{-1}_{\rm MFPI}|x^{(k+1)}|,\end{cases}$ (1.10)
where $Q_{\rm MFPI}\in\mathbb{R}^{n\times n}$ is a nonsingular matrix and $\tau_{\rm MFPI}>0$. Generally, NSOR (1.9) and MFPI (1.10) are different from each other.
Though the RMS iteration (1.3) contains many existing iterations as its special cases, after further investigation, it can not include the general NSOR iteration (1.9) and the general MFPI (1.10) (see the next section for more details). This motivates us to generalize the RMS iteration (1.3) such that, at least, the generalization includes the general NSOR iteration (1.9) and the general MFPI (1.10) as its special cases. To this end, the momentum acceleration technique [30, 23, 2, 22] enters our vision. Concretely, by incorporating the momentum acceleration technique and a new matrix splitting into the RMS iteration (1.3), a generalization of the RMS (GRMS) iteration is developed for solving the GAVE (1.1).
The rest of this paper is organized as follows. In Section 2, we establish the GRMS iterative method for solving the GAVEs (1.1). In Section 3, the convergence theories of the GRMS method are presented in detail. In Section 4, the feasibility and efficiency of the GRMS method are verified by numerical experiments. Finally, the conclusion of this paper is given in Section 5.
Notation. $\mathbb{N}^{+}$ is the set of all positive integer. $\mathbb{R}^{n\times n}$ is the set of all $n\times n$ real matrices, $\mathbb{R}^{n}=\mathbb{R}^{n\times 1}$ and $\mathbb{R}^{+}$ represents the set of the positive real number. $|U|\in\mathbb{R}^{m\times n}$ denote the componentwise absolute value of matrix $U$. $I$ denotes the identity matrix with suitable dimension. $\|U\|$ denotes the $2$-norm of $U$ which is define by the formula $\|U\|=\max\{\|Ux\|:x\in\mathbb{R}^{n},\|x\|=1\}$, where $\|x\|$ is the $2$-norm. For any matrices $U=(u_{ij})$ and $V=(v_{ij})\in\mathbb{R}^{n\times n}$, $U\leq V$ means that $u_{ij}\leq v_{ij}$ for $i,j=1,2,\cdots,n$.
@@ 2 The GRMS iterative method
Let $|x|=Qy$ with $Q\in\mathbb{R}^{n\times n}$, then the GAVEs (1.1) can be transformed into the following two-by-two block nonlinear equation
$\begin{cases}Ax-BQy=b,\\ Qy-|x|=0.\end{cases}$ (2.1)
By splitting $A$ and $Q$ as $A=M-N$ and $Q=Q_{1}-Q_{2}$ with $M$ and $Q_{1}\in\mathbb{R}^{n\times n}$ being invertible. Then the system (2.1) can be rewritten as
$\begin{cases}Mx=Nx+BQy+b,\\ Q_{1}y=Q_{1}y-\theta Q_{1}y+\theta Q_{2}y+\theta|x|+Hx-Hx,\end{cases}$ (2.2)
where $\theta>0$ is a scaled parameter and $H\in\mathbb{R}^{n\times n}$ is an any given matrix. Based on (2.2), the following Algorithm 2.1 can be developed for solving the GAVEs (1.1).
Algorithm 2.1.
Assume that $A,B\in\mathbb{R}^{n\times n}$ and $Q\in\mathbb{R}^{n\times n}$ are given. Let $A=M-N$ and $Q=Q_{1}-Q_{2}$ with $M$ and $Q_{1}$ being invertible. Given initial vectors $x^{(0)}\in\mathbb{R}^{n}$ and $y^{(0)}\in\mathbb{R}^{n}$, for $k=0,1,2,\cdots$ until the iteration sequence $\{(x^{(k)},y^{(k)})\}^{\infty}_{k=0}$ is convergent, compute
$\begin{cases}x^{(k+1)}=M^{-1}Nx^{(k)}+M^{-1}BQy^{(k)}+M^{-1}b,\\ y^{(k+1)}=(1-\theta)y^{(k)}+\theta{Q_{1}}^{-1}Q_{2}y^{(k)}+\theta{Q_{1}}^{-1}|x^{(k+1)}|+{Q_{1}}^{-1}Hx^{(k+1)}-{Q_{1}}^{-1}Hx^{(k)},\end{cases}$ (2.3)
@@
Comparing with the RMS iteration (1.3), the iteration (2.3) involves a new matrix $Q$ as well as its splitting and the momentum acceleration term ${Q_{1}}^{-1}H(x^{(k+1)}-x^{(k)})$ whenever $H\neq 0$, which make the iteration (2.3) more general than the RMS iteration (1.3). Indeed, if $M=M_{\rm RMS},Q=I$, $Q_{1}=\frac{\theta}{\tau_{\rm RMS}}I$, and $H=0$, the iteration (2.3) turns into the RMS iteration (1.3). That is why the iteration (2.3) is called as the GRMS iteration. This also means that the GRMS iteration (2.3) includes the existing iterative schemes contained in the RMS iteration (1.3) as its special cases. Moreover, the GRMS iteration (2.3) contains some iterative schemes which are not generally contained by the RMS iteration (1.3) as its special cases, examples are as follows.
@@
If $B=Q=I$, $M=\frac{1}{\alpha_{\rm MSOR}\omega_{\rm MSOR}}A$, $Q_{1}=\frac{\theta(2-\omega_{\rm MSOR})}{2\omega_{\rm MSOR}}I$ and $H=\frac{\theta(1-\alpha_{\rm MSOR}\omega_{\rm MSOR})}{\alpha_{\rm MSOR}\omega_{\rm MSOR}}A$ ($\alpha_{\rm MSOR},\omega_{\rm MSOR}>0$ and $\omega_{\rm MSOR}\neq 2$), the GRMS iteration (2.3) reduces to
$\begin{cases}x^{(k+1)}=(1-\alpha_{\rm MSOR}\omega_{\rm MSOR})x^{(k)}+\alpha_{\rm MSOR}\omega_{\rm MSOR}A^{-1}(y^{(k)}+b),\\ y^{(k+1)}=y^{(k)}+\frac{2\omega_{\rm MSOR}}{2-\omega_{\rm MSOR}}(-Ax^{(k+1)}+|x^{(k+1)}|+b),\end{cases}$ (2.4)
@@
which is the modified SOR-like (MSOR) iteration proposed in [7] in the context of AVEs (1.2) associated with second order cones. Particularly, if $\alpha_{\rm MSOR}=\frac{2\sigma_{\rm NSOR}+\omega_{\rm NSOR}}{2}$ and $\omega_{\rm MSOR}=\frac{2\omega_{\rm NSOR}}{2\sigma_{\rm NSOR}+\omega_{\rm NSOR}}$, the iteration (2.4) reduces to the NSOR iteration (1.9). However, the RMS iteration (1.3) can only contain the NSOR iteration (1.9) with $\omega_{\rm NSOR}=1$.
@@
If $B=I$, $M=\frac{1}{\alpha_{\rm MGSOR}}A$, $Q=Q_{\rm MGSOR}$, $Q_{1}=\frac{\theta}{\beta_{\rm MGSOR}}Q_{\rm MGSOR}$ and $H=0$ , then the GRMS iteration (2.3) turns into the modified generalized SOR-like (MGSOR) iteration [31]
$\begin{cases}x^{(k+1)}=(1-\alpha_{\rm MGSOR})x^{(k)}+\alpha_{\rm MGSOR}A^{-1}(Q_{\rm MGSOR}y^{(k)}+b),\\ y^{(k+1)}=(1-\beta_{\rm MGSOR})y^{(k)}+\beta_{\rm MGSOR}Q_{\rm MGSOR}^{-1}|x^{(k+1)}|,\end{cases}$ (2.5)
@@
where $\alpha_{\rm MGSOR}$ and $\beta_{\rm MGSOR}$ are positive constants and $Q_{\rm MGSOR}\in\mathbb{R}^{n\times n}$ is a nonsingular matrix. Specially, if $\alpha_{\rm MGSOR}=\beta_{\rm MGSOR}=\omega_{\rm SOR}$ and $Q_{\rm MGSOR}=I$, the MGSOR iteration (2.5) reduces to the SOR-like iteration (1.7). If $\alpha_{\rm MGSOR}=1$, $\beta_{\rm MGSOR}=\tau_{\rm MFPI}$ and $Q_{\rm MGSOR}=Q_{\rm MFPI}$, the MGSOR iteration (2.5) reduces to the MFPI iteration (1.10) [31]. However, the RMS iteration (1.3) can not contain the MGSOR iteration (2.5) with $Q_{\rm MGSOR}\neq I$ and the MFPI iteration (1.10) with $Q_{\rm MFPI}\neq I$.
@@
If $M=M_{\rm MAMS}+\Omega_{\rm MAMS}$ with $A=M_{\rm MAMS}-N_{\rm MAMS}$ and $\Omega_{\rm MAMS}$ being a positive diagonal matrix, $B=Q=I$, $Q_{1}=\theta I$ and $H=\theta\beta_{\rm MAMS}I$ ($\beta_{\rm MAMS}$ is a real number called momentum factor), then the GRMS iteration (2.3) becomes the momentum acceleration-based matrix splitting (MAMS) iteration [32]
$x^{(k+2)}=(M_{\rm MAMS}+\Omega_{\rm MAMS})^{-1}\left[(N_{{\rm MAMS}}+\Omega_{\rm MAMS})x^{(k+1)}+|x^{(k+1)}|+\beta_{\rm MAMS}(x^{(k+1)}-x^{(k)})+b\right]$ (2.6)
@@
with $x^{(1)}=(M_{\rm MAMS}+\Omega_{\rm MAMS})^{-1}[(N_{{\rm MAMS}}+\Omega_{\rm MAMS})x^{(0)}+y^{(0)}+b]$, where $x^{(0)}$ is used for both iterations and $y^{(0)}$ is used in the GRMS iteration. In other words, with suitable initial vectors, the GRMS iteration (2.3) is the same with the MAMS iteration (2.6) for solving the AVEs (1.2). However, it seems impossible to establish such a relationship between the RMS iteration (1.3) and the MAMS iteration (2.6) with $\beta_{\rm MAMS}\neq 0$.
@@
In conclusion, the GRMS iteration (2.3) is more general than the RMS iteration (1.3). Before ending this section, we will summarize more specific versions of the GRMS iteration (2.3) for solving the GAVEs (1.1).

### v2

@@
Abstract
By incorporating a new matrix splitting and the momentum acceleration into the relaxed-based matrix splitting (RMS) method [27], a generalization of the RMS (GRMS) iterative method for solving the generalized absolute value equations (GAVEs) is proposed. On the one hand, unlike some existing methods, by using the Cauchy’s convergence principle we give some sufficient conditions for the existence and uniqueness of the solution to GAVEs and prove that our method can converge to the unique solution of GAVEs. On the other hand, we obtain a few new and weaker convergence conditions for some existing methods. Moreover, we establish comparison theorems between GRMS method and some existing methods. Preliminary numerical experiments show that the proposed method is efficient.
keywords:
Generalized absolute value equations, Matrix splitting, Relaxation technique, Momentum acceleration, Convergence
@@ 1 Introduction
We focus on seeking a solution of the following system of generalized absolute value equations (GAVEs)
$Ax-B|x|-b=0,$ (1)
where $A,B\in\mathbb{R}^{n\times n}$ and $b\in\mathbb{R}^{n}$ are given, and $x\in\mathbb{R}^{n}$ is unknown. Here, $|x|\in\mathbb{R}^{n}$ denotes the vector whose $i$th component is $|x_{i}|$, i.e., the absolute value of $x_{i}$. When $B=I$, GAVEs (1) reduces to the system of absolute value equations (AVEs)
$Ax-|x|-b=0.$ (2)
To the best of our knowledge, GAVEs (1) is formally introduced by Rohn in [25]. Over the past two decades, GAVEs (1) and AVEs (2) have attracted more and more attention in the optimization community and a great deal of research have been done; see, e.g., [17, 7, 9, 13, 18, 19, 20, 4, 12, 1, 21, 5]. Particularly, solving the general GAVEs (1) is NP-hard [18].
Within the numerical algorithms for solving GAVEs (1), some are characterized by using the matrix splitting technique or the relaxation technique. Recently, by using the matrix splitting technique and the relaxation technique, Song and Song [27] present the relaxation-based matrix splitting (RMS) iterative method (see Algorithm 1 below) for solving GAVEs (1).
Algorithm 1
Assume that $A=M_{\rm RMS}-N_{\rm RMS}$ with $M_{\rm RMS}$ being nonsingular. Given any initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$, for $k=0,1,2,\ldots$ until the iterative sequence $\{(x^{(k)},y^{(k)})\}^{\infty}_{k=0}$ is convergent, compute
$\begin{cases}x^{(k+1)}=M_{\rm RMS}^{-1}(N_{\rm RMS}x^{(k)}+By^{(k)}+b),\\ y^{(k+1)}=(1-\tau_{\rm RMS})y^{(k)}+\tau_{\rm RMS}|x^{(k+1)}|,\end{cases}$ (3)
@@
The RMS iterative scheme (3) includes many existing iterative schemes as its special cases [27]. For instance,
@@
    Let $A=M_{\rm NMS}-N_{\rm NMS}$ and $\Omega_{\rm NMS}$ be a given matrix such that $M_{\rm NMS}+\Omega_{\rm NMS}$ is invertible. If $\tau_{\rm RMS}=1$ and $M_{\rm RMS}=M_{\rm NMS}+\Omega_{\rm NMS}$, the RMS iteration (3) with $y^{(0)}=|x^{(0)}|$ turns into the Newton-based matrix splitting (NMS) iteration [34]
    $x^{(k+1)}=(M_{\rm NMS}+\Omega_{\rm NMS})^{-1}[(N_{\rm NMS}+\Omega_{\rm NMS})x^{(k)}+B|x^{(k)}|+b].$ (4)
    The NMS iteration (4) reduces to the modified Newton-type (MN) iteration [28] when $M_{\rm NMS}=A$ and $\Omega_{\rm NMS}$ is a positive semi-definite matrix. Moreover, if $M_{\rm NMS}=A$ and $\Omega_{\rm NMS}=0$, the NMS iteration (4) reduces to the Picard iteration [26]
    $x^{(k+1)}=A^{-1}(B|x^{(k)}|+b).$ (5)
    If $M_{\rm NMS}=\dfrac{1}{2}(A+\Omega_{\rm SSMN})$ and $\Omega_{\rm NMS}=0$, then the NMS iteration (4) turns into the following shift-splitting modified Newton-type (SSMN) iteration [15]
    $x^{(k+1)}=(A+\Omega_{\rm SSMN})^{-1}\left[(\Omega_{\rm SSMN}-A)x^{(k)}+2B|x^{(k)}|+2b\right],$ (6)
    where $\Omega_{\rm SSMN}$ is a positive semi-definite matrix such that $A+\Omega_{\rm SSMN}$ is invertible.
@@
    If $B=I,M_{\rm RMS}=\dfrac{1}{\omega_{\rm SOR}}A$ and $\tau_{\rm RMS}=\omega_{\rm SOR}>0$, the RMS iteration (3) becomes the SOR-like iteration [10]
    $\begin{cases}x^{(k+1)}=(1-\omega_{\rm SOR})x^{(k)}+\omega_{\rm SOR}A^{-1}\left(y^{(k)}+b\right),\\ y^{(k+1)}=(1-\omega_{\rm SOR})y^{(k)}+\omega_{\rm SOR}|x^{(k+1)}|.\end{cases}$ (7)
@@
    Throughout this paper, let $A=D-L-U$ with $D$, $-L$ and $-U$ being the diagonal part, the strictly lower-triangular part and the strictly upper-triangular part of $A$, respectively. If $B=I$, $M_{\rm RMS}=\dfrac{1}{\omega}D-L$ and $\tau_{\rm RMS}=\omega>0$, the RMS iteration (3) turns into the modified SOR-like iteration [14, Equation (11)].
@@
    If $B=I,M_{\rm RMS}=A,\tau_{\rm RMS}=\tau_{\rm FPI}>0$, the RMS iteration (3) is reduced to the fixed point iteration (FPI) [11]
    $\begin{cases}x^{(k+1)}=A^{-1}(y^{(k)}+b),\\ y^{(k+1)}=\left(1-\tau_{\rm FPI}\right)y^{(k)}+\tau_{\rm FPI}|x^{(k+1)}|,\end{cases}$ (8)
@@
    $\begin{cases}x^{(k+1)}=x^{(k)}+\omega_{\rm NSOR}A^{-1}(b-Ax^{(k)}+y^{(k)}),\\ y^{(k+1)}=y^{(k)}+\dfrac{\omega_{\rm NSOR}}{\sigma_{\rm NSOR}}(b+|x^{(k+1)}|-Ax^{(k+1)}),\end{cases}$ (9)
    where $\omega_{\rm NSOR},\sigma_{\rm NSOR}>0$. Indeed, if $\omega_{\rm NSOR}=1$ and $\sigma_{\rm NSOR}=\dfrac{1}{\tau_{\rm FPI}}$, the NSOR iteration (9) reduces to (8). In addition, the RMS iteration (3) is equivalent to the NSOR iteration (9) if $B=I$, $M_{\rm RMS}=A$, $\tau_{\rm RMS}=\frac{1}{\sigma_{\rm NSOR}}$ and $\omega_{\rm NSOR}=1$. Under these conditions, both of them can be seem as the FPI iteration.
Before going ahead, we should point out that the FPI (8) is also a special case of the modified FPI (MFPI) [30]
$\begin{cases}x^{(k+1)}=A^{-1}(Q_{\rm MFPI}y^{(k)}+b),\\ y^{(k+1)}=(1-\tau_{\rm MFPI})y^{(k)}+\tau_{\rm MFPI}Q^{-1}_{\rm MFPI}|x^{(k+1)}|,\end{cases}$ (10)
where $Q_{\rm MFPI}\in\mathbb{R}^{n\times n}$ is a nonsingular matrix and $\tau_{\rm MFPI}>0$. Generally, NSOR (9) and MFPI (10) are different from each other.
Though the RMS iteration (3) contains many existing iterations as its special cases, after further investigation, it cannot include the general NSOR iteration (9) and the general MFPI (10) (see the next section for more details). This motivates us to generalize the RMS iteration (3) such that, at least, the generalization includes the general NSOR iteration (9) and the general MFPI (10) as its special cases. To this end, the momentum acceleration technique [31, 24, 3, 23] enters our vision. Concretely, by incorporating the momentum acceleration technique and a new matrix splitting into the RMS iteration (3), a generalization of the RMS (GRMS) iteration is developed for solving GAVEs (1).
The rest of this paper is organized as follows. In Section 2, we establish the GRMS iterative method for solving GAVEs (1). The convergence theories of the GRMS method are presented in Section 3. In Section 4, comparison theorems between GRMS method and some existing methods are proposed. In Section 5, the feasibility and efficiency of the GRMS method are verified by numerical experiments. Finally, the conclusion of this paper is given in Section 6.
Notation. $\mathbb{N}^{+}$ is the set of all positive integer. $\mathbb{R}^{n\times n}$ is the set of all $n\times n$ real matrices, $\mathbb{R}^{n}=\mathbb{R}^{n\times 1}$ and $\mathbb{R}^{+}$ represents the set of the positive real number. $|U|\in\mathbb{R}^{n\times n}$ denote the componentwise absolute value of matrix $U$. $I$ denotes the identity matrix with suitable dimension. For $U\in\mathbb{R}^{n\times n}$, $\rho(U)$ denotes the spectral radius of $U$, $\|U\|$ denotes the $2$-norm of $U$ which is defined by the formula $\|U\|=\max\{\|Ux\|:x\in\mathbb{R}^{n},\|x\|=1\}$, where $\|x\|$ is the $2$-norm of the vector $x$. For any matrices $U=(u_{ij})$ and $V=(v_{ij})\in\mathbb{R}^{n\times n}$, $U\leq V$ ($U<V$) means that $u_{ij}\leq v_{ij}$ ($u_{ij}<v_{ij}$) for $i,j=1,2,\ldots,n$.
@@ 2 The GRMS iterative method
Let $|x|=Qy$ with $Q\in\mathbb{R}^{n\times n}$. Then GAVEs (1) can be transformed into the following two-by-two block nonlinear equations
$\begin{cases}Ax-BQy=b,\\ Qy-|x|=0.\end{cases}$ (11)
By splitting $A$ and $Q$ as $A=M-N$ and $Q=Q_{1}-Q_{2}$ with $M$ and $Q_{1}\in\mathbb{R}^{n\times n}$ being invertible. Then the system (11) can be rewritten as
$\begin{cases}Mx=Nx+BQy+b,\\ Q_{1}y=Q_{1}y-\theta Q_{1}y+\theta Q_{2}y+\theta|x|+Hx-Hx,\end{cases}$ (12)
where $\theta>0$ is a scaled parameter and $H\in\mathbb{R}^{n\times n}$ is a given matrix. Based on (12), the following Algorithm 2 (which is called as the GRMS iteration method) can be developed for solving GAVEs (1).
Algorithm 2
Assume that $A,B\in\mathbb{R}^{n\times n}$ and $Q\in\mathbb{R}^{n\times n}$ are given. Let $A=M-N$ and $Q=Q_{1}-Q_{2}$ with $M$ and $Q_{1}$ being invertible. Given initial vectors $x^{(0)}\in\mathbb{R}^{n}$ and $y^{(0)}\in\mathbb{R}^{n}$, for $k=0,1,2,\ldots$ until the iteration sequence $\{(x^{(k)},y^{(k)})\}^{\infty}_{k=0}$ is convergent, compute
$\begin{cases}x^{(k+1)}=M^{-1}Nx^{(k)}+M^{-1}BQy^{(k)}+M^{-1}b,\\ y^{(k+1)}=(1-\theta)y^{(k)}+\theta{Q_{1}}^{-1}Q_{2}y^{(k)}+\theta{Q_{1}}^{-1}|x^{(k+1)}|+{Q_{1}}^{-1}Hx^{(k+1)}-{Q_{1}}^{-1}Hx^{(k)},\end{cases}$ (13)
@@
Comparing with the RMS iteration (3), the GRMS iteration (13) involves a new matrix $Q$ as well as its splitting and the momentum acceleration term ${Q_{1}}^{-1}H(x^{(k+1)}-x^{(k)})$ whenever $H\neq 0$, which make the GRMS iteration (13) more general than the RMS iteration (3). Indeed, if $M=M_{\rm RMS},Q=I$, $Q_{1}=\frac{\theta}{\tau_{\rm RMS}}I$, and $H=0$, the GRMS iteration (13) turns into the RMS iteration (3). This also means that the GRMS iteration (13) includes the existing iterative schemes contained in the RMS iteration (3) as its special cases. Moreover, the GRMS iteration (13) contains some iterative schemes which are not generally contained by the RMS iteration (3), examples are as follows.
@@
If $B=Q=I$, $M=\frac{1}{\alpha_{\rm MSOR}\omega_{\rm MSOR}}A$, $Q_{1}=\frac{\theta(2-\omega_{\rm MSOR})}{2\omega_{\rm MSOR}}I$ and $H=\frac{\theta(1-\alpha_{\rm MSOR}\omega_{\rm MSOR})}{\alpha_{\rm MSOR}\omega_{\rm MSOR}}A$ ($\alpha_{\rm MSOR},\omega_{\rm MSOR}>0$ and $\omega_{\rm MSOR}\neq 2$), the GRMS iteration (13) reduces to
$\begin{cases}x^{(k+1)}=(1-\alpha_{\rm MSOR}\omega_{\rm MSOR})x^{(k)}+\alpha_{\rm MSOR}\omega_{\rm MSOR}A^{-1}(y^{(k)}+b),\\ y^{(k+1)}=y^{(k)}+\frac{2\omega_{\rm MSOR}}{2-\omega_{\rm MSOR}}(-Ax^{(k+1)}+|x^{(k+1)}|+b),\end{cases}$ (14)
@@
which is the modified SOR-like (MSOR) iteration proposed in [8] in the context of AVEs (2) associated with second order cones. Particularly, if $\alpha_{\rm MSOR}=\frac{2\sigma_{\rm NSOR}+\omega_{\rm NSOR}}{2}$ and $\omega_{\rm MSOR}=\frac{2\omega_{\rm NSOR}}{2\sigma_{\rm NSOR}+\omega_{\rm NSOR}}$, the iteration (14) reduces to the NSOR iteration (9). However, the RMS iteration (3) can only contain the NSOR iteration (9) with $\omega_{\rm NSOR}=1$.
@@
If $B=I$, $M=\frac{1}{\alpha_{\rm MGSOR}}A$, $Q=Q_{\rm MGSOR}$, $Q_{1}=\frac{\theta}{\beta_{\rm MGSOR}}Q_{\rm MGSOR}$ and $H=0$ , then the GRMS iteration (13) turns into the modified generalized SOR-like (MGSOR) iteration [32]
$\begin{cases}x^{(k+1)}=(1-\alpha_{\rm MGSOR})x^{(k)}+\alpha_{\rm MGSOR}A^{-1}(Q_{\rm MGSOR}y^{(k)}+b),\\ y^{(k+1)}=(1-\beta_{\rm MGSOR})y^{(k)}+\beta_{\rm MGSOR}Q_{\rm MGSOR}^{-1}|x^{(k+1)}|,\end{cases}$ (15)
@@
where $\alpha_{\rm MGSOR}$ and $\beta_{\rm MGSOR}$ are positive constants and $Q_{\rm MGSOR}$ is a nonsingular matrix. Specially, if $\alpha_{\rm MGSOR}=\beta_{\rm MGSOR}=\omega_{\rm SOR}$ and $Q_{\rm MGSOR}=I$, the MGSOR iteration (15) reduces to the SOR-like iteration (7). If $\alpha_{\rm MGSOR}=1$, $\beta_{\rm MGSOR}=\tau_{\rm MFPI}$ and $Q_{\rm MGSOR}=Q_{\rm MFPI}$, the MGSOR iteration (15) reduces to the MFPI iteration (10) [32]. However, the RMS iteration (3) cannot contain the MGSOR iteration (15) with $Q_{\rm MGSOR}\neq I$ and the MFPI iteration (10) with $Q_{\rm MFPI}\neq I$.
@@
If $M=M_{\rm MAMS}+\Omega_{\rm MAMS}$ with $A=M_{\rm MAMS}-N_{\rm MAMS}$ and $\Omega_{\rm MAMS}$ being a positive diagonal matrix, $B=Q=I$, $Q_{1}=\theta I$ and $H=\theta\beta_{\rm MAMS}I$ ($\beta_{\rm MAMS}$ is a real number called momentum factor), the GRMS iteration (13) becomes the momentum acceleration-based matrix splitting (MAMS) iteration [33]
$x^{(k+2)}=(M_{\rm MAMS}+\Omega_{\rm MAMS})^{-1}\left[(N_{{\rm MAMS}}+\Omega_{\rm MAMS})x^{(k+1)}+|x^{(k+1)}|+\beta_{\rm MAMS}(x^{(k+1)}-x^{(k)})+b\right]$ (16)
@@
with $x^{(1)}=(M_{\rm MAMS}+\Omega_{\rm MAMS})^{-1}[(N_{{\rm MAMS}}+\Omega_{\rm MAMS})x^{(0)}+y^{(0)}+b]$, where $x^{(0)}$ is used for both iterations and $y^{(0)}$ is used in the GRMS iteration. In other words, with suitable initial vectors, the GRMS iteration (13) is the same with the MAMS iteration (16) for solving AVEs (2). However, it seems impossible to establish such a relationship between the RMS iteration (3) and the MAMS iteration (16) with $\beta_{\rm MAMS}\neq 0$.
@@
In conclusion, the GRMS iteration (13) is more general than the RMS iteration (3). Before ending this section, we will summarize more specific versions of the GRMS iteration (13) for solving GAVEs (1).